### v3

# Ravensburg

Abstract
GenL is a flexible program that can be used to simulate and/or fit x-ray diffraction data from epitaxial thin films exhibiting Laue oscillations. It utilizes differential evolution within a genetic algorithm for the fitting of data and is based on the kinematic theory of diffraction. Effects of polarization, absorption, the Lorentz factor, as well as instrumental resolution and lattice vibrations are taken into account. Useful parameters that can be extracted after fitting include: atomic interplanar spacings, number of coherently scattering atomic planes, strain profiles along the film thickness, and crystal roughness. The program has been developed in MATLAB and employs a graphical user interface. The deployment strategy is twofold whereby the software can either be obtained in source code form and executed within the MATLAB environment, or as a pre-compiled binary for those who prefer not to run it within MATLAB. Finally, GenL can be easily extended to simulate multilayered film systems, superlattices, and films with atomic steps. The program is released under the GNU General Public Licence.
Supply a synopsis of the paper for inclusion in the Table of Contents.
1 Introduction and motivation
High crystal quality epitaxial thin films are intriguing model systems. Low defect density and minor crystal misorientation allow for a direct and easy comparison to theoretical predictions [Neugebauer2013]. To experimentally verify the crystal quality of epitaxial thin films, the presence of Laue oscillations in x-ray diffraction (XRD) patterns is regarded as a key indication of high crystallinity [Miller2022]. Laue oscillations are only observable if the coherence length of the epitaxial layer is of the order of the film thickness and the latter is smaller than the longitudinal coherence length of the x-ray photons. The occurrence of Laue oscillations is often used to qualitatively characterize samples, associating their presence with high crystallinity and low defect density [Krauss2008, Shu2020, Song2020, Kamigaki1990, DeTheresa2007], thickness uniformity [Naito2001, Xu2017, Hu2013], smooth and flat interface layering [Abe2002, Xu2017, Miyadera2021, Radu2006], and film homogeneity [Forst2015]. Nevertheless, a quantitative analysis of the Laue oscillations, modeling their angular spacing, intensity decay, and asymmetry around their respective Bragg peak can yield additional structural insights [Fullerton1992, Stierle1993, Komar2017, Miller2022].
For instance, the characteristic period of the Laue oscillations is related to the number of coherently scattering layers, and hence, to the coherent thickness of the layer. However, the thickness determined via the separation of Laue oscillations can deviate from the thickness determined by the spacing of the Kiessig fringes found in reflectivity measurements [Kiessig1931]. A possible origin of this may be the presence of excess material at an interface [Miller2022] with a similar electron density as the film but with a different crystal structure. This region adds to the total film thickness determined via the spacing of the Kiessig fringes but it would not affect the thickness determined via the Laue oscillations as the mismatch in crystal structure terminates the crystal coherence length of the film. The intensity decay of the Laue oscillations can be understood in terms of the roughness of the layer, i.e., the distribution of out-of-plane thickness that contributes to coherent scattering. Moreover, an asymmetry of the Laue oscillation intensity on either side of the Bragg peak is frequently observed [Miller2022] and often associated with out-of-plane strain [Vartanyants2000, Robinson2001, Kastle2002, Komar2017], or other factors [Miller2022]. Numerous powerful open-source program codes exist for simulating and fitting of diffraction patterns including Laue oscillations, e.g., SUPREX [Fullerton1992], GenX [Bjorck2007, Glavic_Bjorck_2022], CADEM [Komar2017], or InteractiveXRDFit [Lichtensteiger2018] alongside with commercially available software, e.g., BedeREFS [Wormington1999] or Leptos [Ulyanenkov2004]. However, the focus of most of them lies mainly on the fitting of superlattice peaks and not on the detailed analysis of the Laue oscillations and their shape. The open-source software SUPREX is designed for superconductor superlattices and written in Fortran and Turbo Pascal [Fullerton1992] with a non-trivial extension to other materials [Komar2017]. GenX is designed to fit x-ray reflectivity patterns [Bjorck2007], being primarily sensitive to electron density profiles along the z-axis. But it can be adjusted to fit high-angle scattering data. Though, reconfiguring the program for these settings is not effortless and detailed analysis of the shape of the oscillations is difficult to implement. CADEM on the other hand, which is written in Matlab, can be used to simulate scattering patterns without a fitting option and without the inclusion of crystal roughness. Finally, InteractiveXRDFit, a Matlab program designed for oxide thin films and heterostructures, utilizes an intermediate approach between the two limits of the kinematic and dynamic x-ray diffraction theories to calculate, but not fit, the diffracted intensity [Lichtensteiger2018]. Besides these, Monte Carlo simulations can be used to reproduce experimental diffraction patterns, however, their application is hindered by their high degree of complexity for modeling realistic sample structures [Komar2017].
Hence, we identified the need for GenL, an open source and easy to adapt Laue oscillation fitting program, which is able to not only simulate but also fit experimental diffraction patterns of epitaxial thin films. In GenL, a representative diffraction pattern, which includes all the aforementioned parameters influencing the shape of the measured oscillations, is simulated and fitted to experimental data. The associated parameters are tuned via a differential evolution algorithm [Storn1997] until a difference minimum of the figure of merit between the data and simulation has been obtained.
In this work, we first present the theoretical background for the calculations of the diffraction patterns used in GenL. We then illustrate how the program can be used to fit experimental data and to analyze the sample layering and crystal structure, with the option to include roughness as well as strain profiles and defects. Finally, we show a few examples of the versatile applications of GenL.
2 Theoretical background
X-ray diffraction can be used to determine the separation between crystallographic planes via Bragg’s law [Bragg1913] :
where is the x-ray wavelength and is the angle at which constructive interference occurs between sets of planes associated with the Miller indices . Equivalently, the conditions for Bragg reflections are fulfilled when the photon’s momentum transfer wave vector coincides with the reciprocal crystal lattice point represented by the vector . When operating a diffractometer in a symmetric /2 configuration, i.e., the detector angle and incident angle are varied in such a way that the incident angle is half of the scattering angle 2, the momentum transfer vector becomes one dimensional and perpendicular to the lattice planes (after alignment for miscut and other factors). Its modulus can then be written as:
The length of is varied by changing the scattering angle (angle between source and detector). Assuming the reciprocal lattice vector is oriented parallel to , and referring to this direction as (with being perpendicular to and in the scattering plane), the intensity diffracted from the crystal can be written in the kinematic approximation as:
where corresponds to the total number of coherently scattering lattice planes within the coherence volume of the x-rays, is the position of the atomic plane in the thin film stack (taking in the center of the first layer above the substrate interface), and is the Debye-Waller factor. The element specific Debye-Waller pre-factor includes the contribution of the time-averaged displacement of atoms due to lattice vibrations. The dispersion corrected and -dependent form factor is given by [Wilson1993, Henke1993]:
along with the anomalous and absorption corrections of Brennan and Cowan [Brennan1992], whereas is commonly expressed in its parameterized form [Waasmaier1995]:
where , , , are the form factor parametrization coefficients of a particular atomic species. For an alloy, is a mole fraction weighted sum of the form factors of the elements in the alloy (with the possibility of having different form factors for different oxidation states).
Several other factors need to be considered [Warren1990, Fullerton1992] if the diffracted intensity is to be compared to the experimentally observed one. The first factor contains the inherent polarization dependence of the scattering of an x-ray with an electron as well as the change in polarization when going through a monochromator. The second factor, , accounts for the absorption of x-rays along the angle dependent path lengths of the incident and diffracted rays. The third factor is the Lorentz factor which takes into account deviations from a perfect instrument and the mosaic nature of the crystal. The experimentally observed intensity is then given by:
Further details on origin and derivation of these factors can be found elsewhere [Buerger1945, Warren1990, Fewster1996, Birkholz2005], but for the convenience of the reader, we give a brief overview of their importance here.
The polarization factor depends on whether a monochromator is used and whether it is located on the incidence side or on the detector side. On the incidence side it takes the form [Azaroff1955]:
| while on the detector side [Yao1982]: | |||
where is the diffraction angle of the monochromator. If no monochromators are being used, then is equal to zero.
The absorption factor corrects for the angle-dependent attenuation of x-rays in the sample. For a single layer measured using a geometry, the absorption factor, corresponding to the effective combined probe and escape depth of the sample, is equal to:
where is the thickness of the film. We assume all extinction effects are negligible. The linear attenuation coefficient can be computed as the product of the atomic density (number of formula units per volume) and the photoabsorption cross-section of the sample, which is included in the form factor correction . The linear attenuation coefficient is then written as:
where is the electron radius, is the number of elements in the alloy, and is the mole fraction of the element .
Finally, the Lorentz factor captures the angular dependent changes of the intensity due to the combination of monochromator imperfections, beam divergence, finite detection resolution, and mosaicity in the sample. Further details can be found elsewhere [Warren1990]. The combined change in intensity can be corrected via:
3 GenL
In GenL, a diffraction pattern is calculated based on these equations. However, to model x-ray scattering intensity measured in a non-ideal diffractometer setup of a non-perfect single crystal, additional intensity contributions are included in the calculation. Instrumental peak broadening is accounted for by a Gaussian convolution. Moreover, the background intensity originating from thermal diffuse scattering, fluorescence radiation, Compton scattering, air scattering, the detector dark current, and coherent diffuse scattering from defects, needs to be included. In the close vicinity of a peak, assuming a linear dependence often allows for a reliable modeling of [Birkholz2005]:
| However, for certain samples, which, e.g., include amorphous layers (for example protective cap layers), a more complex polynomial modeling might be needed, such that: | |||
where , , and are constants. Both options are available to choose from in GenL. Furthermore, based on the user’s choice, the intensity calculation in GenL can include an additional intensity originating from a single crystalline substrate. Due to the large extension of the substrate compared to the film, the substrate Bragg peak is modeled with a Lorentzian function following:
where is the substrate peak intensity, is the half width at half maximum of the peak and is the expected peak position in . However, for the case that the Bragg peak positions from film and substrate lie close to each other, scattering from film and substrate will interfere, which can cause an asymmetry in the Laue oscillations as well as an intensity shift of the Bragg peak position [Robinson1988]. An accurate calculation of the scattering pattern hence requires to add the scattering amplitudes from film and substrate prior to squaring. It is possible to manually adapt the GenL command line version to simulate a semi-infinite stack of atoms from a substrate, a substrate/film interface, and the atomic stack of the epitaxial layer on top, based on the work presented elsewhere [Als-Nielsen.ch3, Als-Nielsen.ch6]. The calculation of scattering amplitudes from a single crystalline substrate will be part of an upcoming version of the program.
In addition, in GenL, out-of-plane strain in the thin film as well as layer roughness can be taken into account. Strain in a single layer is modeled by displacing the atoms in the calculated one-dimensional stack according to a certain strain profile. The displacement of the atom from its position in an unstrained lattice [Fullerton1992] is given by:
where denotes the out-of-plane distance between the atomic planes for the unstrained case and is a fitting parameter. In the program we consider strain originating from the substrate/film interface () and strain originating from an additional film/capping interface (). The latter is not relevant for sole single layers without capping. The respective induced strain can either be tensile or compressive, resulting in the addition or subtraction of to/from the position in an unstrained lattice, respectively. In GenL it is possible to employ either or both of the strain profiles. The layer roughness is included in GenL employing:
where the set are all integers in a interval around , i.e., [Fewster1996]. In the command line version of GenL, any other model for strain and layer roughness can be manually included, e.g., strain due to charge density waves [Singer2016] or roughness based on binomial fluctuations [Miceli1992].
In order to then match the theoretical x-ray scattering pattern calculated in GenL and the measured intensity as best as possible, 13 parameters can be fitted in the program:
the out-of-plane interplanar spacing
the number of coherently scattering planes
resolution and scale factor for the Gaussian convolution
parameters modeling the background intensity , , and
parameters modeling a substrate peak , ,
the strain parameters and
the roughness parameter
The fitting in GenL utilizes differential evolution within a genetic algorithm [Storn1997] like the one used in GenX [Bjorck2007, Glavic_Bjorck_2022]. The algorithm first generates parameter vectors. The population of parameter vectors (parent population) is then changed based on a certain process to create a new population with a better fit compared to the previous one until the stopping criterion – in GenL the maximum number of iterations – is fulfilled. Further details on the used algorithm as well as on the population generation can be found elsewhere [Bjorck2007]. The genetic algorithm can be tuned in the command line version of GenL in terms of the population size and crossover factor.
GenL is similar to GenX as it also offers a graphical user interface (GUI), both are open-source and can be customized individually for specific applications. Furthermore, both have embedded libraries/interfaces with databases to fetch material and scattering parameters. However, in GenX the layering in a thin film stack is simulated, whereas GenL simulates the atomic ordering. It provides deeper insights into the atomic structures and crystal properties, which GenX fits cannot do. In contrast to GenX, GenL is not applicable to all kinds of layers, e.g., amorphous layers.
An example of such a GenL fit of the scattering pattern of an epitaxial vanadium thin film is displayed in Fig. 1. Based on XRR measurements, the thickness of the V layer is 105 Å. The fit takes the tensile out-of-plane strain due to the negative lattice mismatch between the vanadium and the MgO () substrate into account. The calculated, exponentially decaying strain profile is shown as an inset. The CADEM simulation for the same data set published by Komar et al. [Komar2017] is displayed as a blue line. Also the CADEM simulation includes a strain profile [Komar2017], which is in qualitative agreement with the strain profile presented here. The magnitude of the exponential decay in out-of-plane lattice parameter as a function of position is larger for the fit with GenL, in particular for depths in the sample corresponding to the first two unit cells. The drastic change in is, hence, calculated to occur only close to the substrate/V interface and might include contributions from interface roughness or atomic steps, which may contribute to an asymmetric decay of the oscillations but are not considered here. The over-estimation of the tetragonal unit cell distortion at the interface is therefore considered to be an artifact. Numeric values of fitted strain profiles should be analyzed carefully and may be considered only a qualitative display if important contributions to the scattering intensity, especially at interfaces, have been neglected in the selected model for a certain sample. The CADEM simulation corresponds to a coherent thickness of 84 Å. Based on the GenL fit, the coherent thickness of this V film is 86.3 Å. Komar et al. [Komar2017] attribute this difference to the by XRR determined thickness to a partly oxidized surface with similar electron density but different crystal structure. We see that we can account for the asymmetry of the profile by including a strain profile combined with crystal roughness.
3.1 Technical details
GenL has been developed in the high-level programming language MATLAB (The MathWorks Inc., Natick, MA, USA). MATLAB is available for numerous operating systems, namely Windows, macOS, and Linux/Unix. The GenL program can either run from the command line after adjusting the input file in a text editor or the GenL GUI can be used. Both methods are equivalent. Since the code is rather compact, an adaptation to other programming languages such as Python, C, C++, or Java is straightforward. However, any alteration of the GUI requires a deeper understanding of the MATLAB programming language. The program was tested on the following MATLAB versions: R2019b, R2020b, R2022b, R2023a, and R2023b. No additional toolbox installations are required. The GenL GUI runs correctly in MATLAB versions published in 2022 or later. In older versions, the graphical display in the app might be distorted. Possible distortions of the graphical display in all versions can be avoided if the program is running on one screen only.
3.2 Using GenL
For running GenL from the command line, the user is required to edit the input file ”fit_laue_oscillations.m”. This input file is displayed as a green trapeze shaped box in the flow chart in Fig. 2. First, file and path of the measurement as well as important measurement parameters need to be specified. Second, material specific parameters (atomic number , density, chemical composition, the Debye-Waller pre-factor), sample parameters (layer thickness, expected Bragg angle/interplanar spacing), and instrument parameters (wavelength, setup, measurement time, amplitude, resolution, and background level) need to be added. For a limited number of elements, the element-specific Debye-Waller pre-factor can be extracted from a database [Sears1991] as well or be put in manually. Furthermore, three features can be included in the fitting process of the program: strain, roughness, and a potentially present peak originating from a single-crystalline substrate can be accounted for by adding a Lorentzian intensity contribution at the respective Bragg angle. If any of these features are included, their respective parameters need to be specified in the code as well. Finally, for all fitting parameters upper and lower bounds are constructed and can be adjusted as necessary. Running the input file will start the fitting process, which is further described in the next section.
3.3 Program organisation
A flow chart of all GenL modules including their function names is displayed in Fig. 2. When GenL is executed, form factors are constructed following equations 4 and 5. In the input file a fitting function is called with start values as well as upper and lower boundaries for all fitting parameters. The fitting function creates a current solution for all parameters, based on which a scattering pattern is calculated according to the equations in section 2 in the evaluation function. The evaluation function compares the theoretical intensity profile and the data input and calculates the goodness of the fit for the current solution using a logarithmic figure of merit. It is straightforward to add other figures of merit to the code. In the fitting function the figures of merit of solutions are compared. The fitting function and the function that compares different solutions were adapted from the work of Storn et al. [Storn1997]. If the current solution is an improvement compared to previous solutions, the fit is plotted and the corresponding parameters are printed out in the command window. A maximum number of iterations is specified by the user in the input file.
3.4 Graphical user interface
GenL can also be used via the GenL GUI. The program can be executed within the MATLAB environment, or through a pre-compiled binary, the latter of which does not require an installation of MATLAB to run. A flow chart of the GUI is displayed in Fig. 3. The GenL GUI consists of three different tabs corresponding to three steps: data import, simulating a sample, and fitting the data. Alternatively, the user can also select an option in the import tab to simulate only without data import or fitting. Following the first path, the user is asked to import the measurement which is to be fitted and prompted to enter the relevant measurement parameters. In the simulate tab, the user supplies all relevant information on the expected crystal structure and layering of the sample as well as on the instrument setup. In the fit tab, the user can proceed to tick a selection of parameters to be fitted. Moreover, a diffraction pattern can at this stage be simulated based on the chosen parameter values and plotted against the imported data. It is possible for the user to display the fit, along with the fitted parameters with their respective lower and upper bounds, the evolution of the figure of merit, and the fitted strain profile.
4 Application examples
In this section we showcase some examples of employing GenL to fit diffraction patterns measured on W and Fe thin films. Details on the thin film growth can be found in the respective publications for the W [Ravensburg2023] and the Fe thin films [Ravensburg2022, Ravensburg2024Fe]. All diffractograms presented in this section were measured in a Bede D1 diffractometer equipped with a Cu x-ray source operated at 35 mA and 50 kV. A circular mask (diameter: 0.005 m) and an incidence and a detector slit (both 0.0005 m) were used. The beam was monochromatized by reducing the CuK and CuK radiation using a Göbel mirror and a 2-bounce-crystal on the incidence side. The x-rays were detected with a Bede EDRc x-ray detector. Details on growth and diffraction study of an Fe/V superlattice with a ratio of 4 monolayers of Fe to 28 monolayers of V in 11 bilayer repetitions, which is presented in comparison to a calculation with GenL, can be found elsewhere [Droulias2017].
4.1 The influence of roughness
Thin W () layers can be epitaxially grown on AlO () substrates with high crystal quality despite a large lattice mismatch of 7.2 % and 19.4 % along W [] and W [] directions [Ravensburg2023]. The diffraction pattern of such a thin film, as shown in Fig. 4, exhibits Laue oscillations around the W () Bragg peak. The Laue oscillations have a high degree of symmetry indicating a small degree of strain in the layer. Fitting the oscillations in GenL yields the fit displayed in blue in Fig. 4. The fit captures the main Bragg peak position and the spacing of the oscillations well. However, the ratio of the height of the oscillations in relation to the main Bragg peak height is overestimated. Taking interface roughness into account in GenL, yields a better match of the relative intensity ratio. The corresponding fit estimating the roughness to be 4 Å is displayed in red. This interface roughness obtained from the GenL fit is in agreement with roughness values at the W/capping interface obtained from Kiessig fringe fitting using GenX [Bjorck2007, Glavic_Bjorck_2022] of the same sample [Ravensburg2023]. The analysis of the GenL fitting employed in [Ravensburg2023] contributed to the understanding that W exhibits a semicoherent interface on AlO () with an immediate strain relaxation taking place directly at the substrate/W interface and serves as an example of the successful application of GenL in studies on epitaxial growth.
4.2 Accounting for overlapping substrate peak intensity
A similar analysis can also be done with the scattering pattern around the W () Bragg peak instead, which is displayed in Fig. 5. However, the sharp peak at around 38 degrees originating from the single crystalline AlO () substrate overlaps in intensity with the Laue oscillations, making an analysis more challenging. This angular overlap in the scattering patterns originating from film and substrate is common for epitaxial growth as it is attributed to the similar out-of-plane lattice parameters in the substrate and the film. In GenL, the substrate peak intensity can be modeled with a Lorentzian profile, which is added to the scattering intensity attributed to the film. An example of a fit including roughness together with this modeled substrate peak intensity is displayed in red. Both fits from Fig. 4 and Fig. 5 for the same sample, yield similar fitted parameters. The deviation of the obtained from both fits is 0.05 %.
4.3 The influence of strain
Besides roughness and substrate peak intensity, GenL is also capable of taking a strain profile, i.e., a change in out-of-plane interplanar spacing over layer thickness into account, similar to [Lichtensteiger2018]. In Fig. 6 the diffraction pattern of a 100 Å thick Fe layer grown on an MgAlO () substrate is displayed. The GenL fit, employing an exponentially decaying strain profile, is shown in blue. Based on the fit, the out-of-plane lattice spacing decreases over Fe layer thickness, in line with a tensile out-of-plane strain induced at the substrate/Fe interface by a lattice mismatch [Ravensburg2022]. Moreover, the shape of this strain profile is in agreement with the observation that thinner Fe films, which are epitaxially grown on MgAlO (), exhibit on average larger out-of-plane lattice spacings in the order of the predicted values in the profile. However, the fit does not fully capture the asymmetric decay of the oscillations on both sides of the Bragg peak. In an experimental study on thin Fe layers (6 to 100 Å) grown on MgAlO () substrates [Ravensburg2024Fe], Fe grows in a tetragonally distorted, body centered tetragonal (bct) crystal structure for Fe layer thicknesses below 3 to 8 ML before the equilibrium body centered cubic (bcc) crystal structure stabilizes on this substrate. GenL can be used to investigate this growth even at the thin film limit. Hence, the diffraction pattern was fitted employing the out-of-plane interplanar distance profile of a bct Fe/bcc Fe bilayer with both bilayer thicknesses and interplanar spacings as fitting parameters. The results are displayed in red in Fig. 6. It is evident that the fit based on the bilayer profile captures the detected scattering intensity better and reproduces the asymmetry present in the intensity of the oscillations around the Fe () Bragg peak. This example demonstrates the full potential of the application of GenL in research work.
4.4 Limitations and possible extensions of the program
Besides the factors already included in GenL, there are others that may also influence the measured diffracted intensity from an epitaxial thin film, e.g., sample alignment, the x-ray beam footprint, crystallographic defects, or mosaicity, all of which are not yet included in the program. Some of these factors will be added in future versions of GenL. Currently, epitaxial thin films composed of more than one element (alloys) can only be modeled via a composition and an average density over the layer, not fully accounting for the correct atom stacking in the unit cell. Including correct atom stacking in alloys requires the computation of a structure factor for the specific crystallographic positions of the respective atoms in the unit cell. Details on such calculations can be found elsewhere [Lichtensteiger2018]. However, the command line version of GenL can be extended to correctly calculate diffraction patterns for these alloys due to the fact that GenL is modeling an effective one-dimensional stack of atoms whose order and positions can be adjusted. Even different/additional layers with multiple repetitions can be included. Hence, GenL can be easily adapted for more complex multilayer stacks composed of different materials and/or different crystal structures, such as superlattice structures. An example of a diffraction pattern from an Fe/V superlattice simulated in GenL, is displayed in Fig. 7 and compared against experimentally measured diffraction data [Droulias2017]. Fitting of superlattice diffraction data is not yet implemented in the current version of GenL, but will be included in future versions. The simulation of superlattice diffraction patterns is, however, straight forward. A version of the GUI, which can simulate superlattice diffraction patterns, is available in the program repository. Moreover, the code versatility also enables the user to conduct a more in-depth analysis by simulating the presence of atomic step terraces. In particular for epitaxial systems where out-of-plane atomic steps in the substrate are incommensurate with the out-of-plane atomic distance in the epitaxially growing film, interference of the scattering intensity from different stacks with vertical mismatch can contribute to an observed asymmetry in Laue oscillations [Fullerton1992, Miller2022], which may be relevant for fitting certain scattering patterns.
5 Conclusions
We developed and present a new versatile program, GenL, for the simulation of diffraction patterns from epitaxial thin films. The combination of this program along other existing x-ray reflectivity fitting programs, such as GenX [Bjorck2007, Glavic_Bjorck_2022], can form a powerful toolbox for rendering a fuller and more detailed structural picture for highly crystalline epitaxial thin films, exploiting x-ray scattering techniques. Finally, the ability to include and fit against experimental data, a series of interface (strain) and structural parameters (atomic steps, superlattices etc) will hopefully be a useful asset for the research community, encouraging a detailed and quantitative analysis of diffraction patterns comprising Laue oscillations.
6 Program distribution
GenL is freely available from the authors under the GNU General Public Licence (GPL). A web page (URL: https://github.com/scatterer/GenL) exists for the distribution of the program. The package contains the Matlab code, an executable of the GUI, example files for fitting and examples for sample simulations if the GUI is used. Comments, feedback, and support can be communicated using the following email addresses: gunnar.palsson@physics.uu.se and vassilios.kapaklis@physics.uu.se.
Acknowledgements: VK would like to acknowledge financial support from the Swedish Research Council (Project No. 2019-03581). GKP acknowledges funding from the Swedish Research Council (Project No. 2018-05200). The authors would like to express their gratitude to Prof. Dr. Gerhard Jakob for supplying the data presented in Fig. 1.
Data availability: The data that support the findings of this study are available from the authors upon reasonable request.